# SARDP_supp

### v1

@@ 4.3. A Real Data Example
We compare the CLE and CLS using a transaction dataset of small- and medium-sized restaurants. The dataset is provided by Shouqianba, which is one of a leading company in China’s mobile payment market (http://www.shouqianbao.com.cn/). This dataset contains information from $N=2024$ fast-food restaurants specializing in snacks in Guangzhou, China. For each restaurant $i$ $(1\leq i\leq N)$, $Y_{i}$ is defined as the transaction volume of the restaurant from April 1st to June 1st, 2024. Subsequently, for each restaurant, we consider three covariates: (1) repeat customers $X_{1}$, which is defined as the percentage of consumers who dined at the restaurant for two or more times in the previous two months; (2) operating hours $X_{2}$, wherein $x_{2i}=1$ indicates that the restaurant operates during both lunch and dinner hours, and $x_{2i}=0$ indicates otherwise; (3) transaction amount per customer $X_{3}$, which is defined as the average amount of each transaction in the previous two months. All the continuous variables are standardized. For privacy protection, the platform has added Gaussian noise to the average transaction amount per customer for restaurants with $\lambda_{x}^{2}=0.2$, and added Gaussian noise to the response $\mathbb{Y}$ with $\lambda^{2}=0.25$.
To analyze the effect of the network structure, we construct a network between restaurants based on the customers who have dined at them. Specifically, if two restaurants $i$ and $j$ have at least five common customers, an edge is defined with $a_{ij}=1$; otherwise, $a_{ij}=0$. Furthermore, the network density (i.e., $\{N(N-1)\}^{-1}\sum_{i_{1},i_{2}}a_{i_{1}i_{2}}$) is 0.31%, which implies that this is a relatively sparse network.
@@
From the results, we can draw the following conclusions. First, it can be observed that the estimation results for both the network effect and covariates using both methods are similar, where the estimated values differ by no larger than 0.04. Second, the network effects estimated by both methods are significantly negative. This is because the network construction is based on customers who have dined at the same restaurants. The more customers who have dined at both restaurants, the stronger the competition between those restaurants. Therefore, the negative network effect can be intuitively explained. Third, the CLS has slightly larger RMSE than that of the CLE. Lastly, the CLS only requires less than one tenth of the computational time of the CLE as it does not involve the calculation of $\Omega^{-1}$.
@@ 5. CONCLUDING REMARKS
This study addressed the issue of privacy protection in the SAR model with noise-added response and covariates. We focused on achieving a consistent estimation of the model parameters and proposed two different estimators: the CLE and the CLS. Both of the estimators provide reliable solutions in privacy-protected scenarios. Theoretical properties have been carefully established.
For future research, we will discuss three potential directions. First, the most common form of the SAR model is discussed here. However, more generalized SAR models are worth exploring, and the properties of estimation after the addition of noise can be investigated. Second, network data is often distributed and stored on different local machines. The challenges of distributed algorithms in privacy-protected scenarios should be addressed. Third, when handling large-scale data, the problem of network sampling often arises. Therefore, addressing the estimation bias caused by network sampling in the presence of noise-added variables is an interesting research problem.

### v2

@@ 4.3. A Real Data Example
We compare the CLE and CLS using a transaction dataset of small- and medium-sized restaurants. The dataset is provided by Shouqianba, which is a leading company in China’s mobile payment market (http://www.shouqianbao.com.cn/). This dataset contains information from $N=2024$ fast-food restaurants specializing in snacks in Guangzhou, China. For each restaurant $i$ $(1\leq i\leq N)$, $Y_{i}$ is defined as the transaction volume of the restaurant from April 1st to June 1st, 2024. Subsequently, for each restaurant, we consider three covariates: (1) repeat customers $X_{1}$, which is defined as the percentage of consumers who dined at the restaurant for two or more times in the previous two months; (2) operating hours $X_{2}$, wherein $x_{2i}=1$ indicates that the restaurant operates during both lunch and dinner hours, and $x_{2i}=0$ indicates otherwise; (3) transaction amount per customer $X_{3}$, which is defined as the average amount of each transaction in the previous two months. All the continuous variables are standardized. For privacy protection, the platform has added Gaussian noise to the average transaction amount per customer for restaurants with $\lambda_{x}^{2}=0.2$, and added Gaussian noise to the response $\mathbb{Y}$ with $\lambda^{2}=0.25$. To analyze the effect of the network structure, we construct a network between restaurants based on the customers who have dined at them. Specifically, if two restaurants $i$ and $j$ have at least five shared customers, an edge is defined with $a_{ij}=1$; otherwise, $a_{ij}=0$. Furthermore, the network density (i.e., $\{N(N-1)\}^{-1}\sum_{i_{1},i_{2}}a_{i_{1}i_{2}}$) is 0.31%, which implies that this is a relatively sparse network.
@@
From the results, we can draw the following conclusions. First, it can be observed that the estimation results for both the network effect and covariates using both methods are similar, where the estimated values differ by no larger than 0.04. Second, the network effects estimated by both methods are significantly negative. This is because the network construction is based on customers who have dined at the same restaurants. The more customers who have dined at both restaurants, the stronger the competition between those restaurants. Thus the negative network effect can be intuitively explained. Third, the CLS has slightly larger RMSE than that of the CLE. Lastly, the CLS only requires less than one tenth of the computational time of the CLE as it does not involve the calculation of $\Omega^{-1}$.
@@ 5. CONCLUDING REMARKS
This study addressed the issue of privacy protection in the SAR model with noise-added response and covariates. We focused on achieving a consistent estimation of the model parameters and proposed two different estimators: the CLE and the CLS. Both of the estimators provide reliable solutions in privacy-protected scenarios. Theoretical properties have been carefully established and model generalizations are discussed.
For future research, we will discuss three potential directions. First, we focused on the classical form of the SAR model and some of its extensions. While these models are well-established, exploring the properties of estimators in dynamic and nonlinear network models with added noise remains an important area for future research. Second, Condition (C5) imposes relatively strict requirements. Investigating how to better protect the privacy of network edges by relaxing these conditions is essential and warrants further study. Finally, examining other methods of privacy protection for the spatial autoregressive model and its extensions presents an intriguing and valuable research direction.